OpenStudy (anonymous):

If I = {integers} and subset P = {positive integers}, what is P'? A:{1, 2, 3, ...} B: {...−3, −2, −1, 0} C:{2, 4, 6, ...} D: { }

OpenStudy (anonymous):

If someone could show me how to do this, it would be greatly appreciated :)

OpenStudy (anonymous):

It's the first one because the positive integers are the "counting numbers" starting with "1" and then "2" and then "3" and so on. That is a subset of all integers, which looks like: { ..., -3, -2, -1, 0, 1, 2, 3, ... }

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

Okay :) Thank you soooooo much!! f f(x) = 2x + 4, find f(3). 10 6 9 4

OpenStudy (anonymous):

The way you would do that is to put "3" where the "x" is on the right side, like so: 2(3) + 4 = 6 + 4 = 10 So, it's the first one.

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

You're always welcome!

OpenStudy (anonymous):